# SKEL-CF: Coarse-to-Fine Biomechanical Skeleton and Surface Mesh Recovery

### v3

@@ 3.2 SKEL-CF
Methods 3DPW [3dpw] Human3.6M [human36m] MOYO [moyo]
MPJPE $\downarrow$ PA-MPJPE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$
HMR2.0 + SKEL fit [SKEL, xia2025hsmr] 81.0 54.4 53.6 34.1 130.5 93.7
HSMR [xia2025hsmr] 81.5 54.8 50.4 32.9 104.5 79.6
SKEL-CF (Ours) 61.5 ($\uparrow$24.5%) 38.7 ($\uparrow$29.3%) 39.0 ($\uparrow$22.6%) 31.2 ($\uparrow$5.4%) 85.0 ($\uparrow$18.6%) 51.4 ($\uparrow$35.4%)
Table 1: Quantitative comparison of skeleton and surface mesh recovery using the SKEL [SKEL] human model on 3DPW [3dpw], Human3.6M [human36m], and MOYO [moyo]. The proposed SKEL-CF delivers consistent accuracy gains over previous approaches, with relative improvements shown in green. denotes results reported by [xia2025hsmr].
@@
Given an input image, the encoder first divides it into visual tokens and, together with camera intrinsic information, extracts contextual features through standard self-attention. Using these features, the encoder produces initial estimates $\boldsymbol{\Theta}_{0}=\{\boldsymbol{\theta}_{0},\boldsymbol{\beta}_{0},\boldsymbol{\pi}_{0}\}$. These initial predictions are then progressively refined through $T$ decoder layers using cross-attention with the encoded visual features as context. The final output of SKEL-CF is $\boldsymbol{\Theta}_{T}=\{\boldsymbol{\theta}_{T},\boldsymbol{\beta}_{T},\boldsymbol{\pi}_{T}\}$.
Given the pose $\boldsymbol{\theta}$ and shape $\boldsymbol{\beta}$, the corresponding $K$ 3D joint positions $\mathbf{J}_{3d}\in\mathbb{R}^{K\times 3}$ can be reconstructed via the forward kinematics function. These 3D joints are then projected onto the full image plane using the camera parameters to obtain 2D joint locations $\mathbf{J}_{2d}\in\mathbb{R}^{K\times 2}$. The resulting 2D joints provide pixel-level supervision, ensuring consistency between the 3D reconstruction and the image evidence. The training supervision includes the pseudo-ground-truth pose $\boldsymbol{\hat{\theta}}$, shape $\boldsymbol{\hat{\beta}}$, 3D joints derived from the pose and shape $\mathbf{\hat{J}}_{3d}$, and 2D keypoint annotations $\mathbf{\hat{J}}_{2d}$. Note that no direct camera supervision $\boldsymbol{\hat{\pi}}$ is provided, the camera parameters are learned solely from 2D keypoint supervision, making the training challenging and weakly supervised.
Learning objectives.
To learn SKEL parameters effectively, we apply multiple supervision signals for accurate 3D human reconstruction, following previous work [xia2025hsmr]. First, we optimize the coarse keypoint loss $\mathcal{L}_{\text{kp}}$, which consists of the $\ell_{1}$ loss for both 3D and 2D joint positions:
@@
Since the proposed dataset 4DHuman-SKEL provides reliable parameter ground truth, we further impose parameter-level supervision for the shape $\boldsymbol{\beta}$ and pose $\boldsymbol{\theta}$, which denotes as $\mathcal{L}_{\beta}$ and $\mathcal{L}_{\theta}$ respectively:
$\mathcal{L}_{\beta}=\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{1}\text{, }\mathcal{L}_{\theta}=\|\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}\|_{1}$ (2)
Finally, to learn target parameter $\boldsymbol{\hat{\Theta}}=\{\boldsymbol{\hat{\theta}},\boldsymbol{\hat{\beta}},\boldsymbol{\hat{\pi}}\}$, we use $\mathcal{L}_{skel}$ which is a combination of the three losses:
$\mathcal{L}_{\text{skel}}=\lambda_{kp}\mathcal{L}_{\text{kp}}+\lambda_{\beta}\mathcal{L}_{\beta}+\lambda_{\theta}\mathcal{L}_{\theta}$ (3)
where $\lambda_{kp}$, $\lambda_{\beta}$ and $\lambda_{\theta}$ are hyper-parameters.
Figure 3: Visual comparison between the proposed SKEL-CF and HSMR [xia2025hsmr]. Our proposed SKEL-CF achieves more precise skeletal estimations (best viewed in zoom). Additional visual results are provided in the supplementary material.
@@
Accurately regressing pose and shape parameters in a single forward pass is challenging, especially for complex articulations such as yoga poses. To mitigate this, SKEL-CF adopts a coarse-to-fine refinement strategy. The encoder first produces an initial estimate of the parameters, denoted as $\boldsymbol{\Theta}_{0}=\{\boldsymbol{\theta}_{0},\boldsymbol{\beta}_{0},\boldsymbol{\pi}_{0}\}$. The decoder then refines these predictions across $L$ layers, where the $i$-th layer outputs $\boldsymbol{\Theta}_{i}=\{\boldsymbol{\theta}_{i},\boldsymbol{\beta}_{i},\boldsymbol{\pi}_{i}\}$. The final layer produces the refined prediction $\boldsymbol{\Theta}_{L}$. Both the encoder output and the final decoder output are optimized using the same objective function $\mathcal{L}_{\text{skel}}$:
@@
To stabilize training, each decoder layer is supervised individually. Specifically, the pose parameters predicted at each layer are optimized with the following refinement loss:
@@
where $L$ denotes the total number of decoder layers.
@@
The overall loss function combines the above components for both the coarse encoder prediction and the refinement decoder:
@@
Following HSMR [xia2025hsmr], we adopt ViTPose-H [xu2022vitpose], pretrained on COCO [coco], as the encoder backbone. ViTPose-H consists of 32 transformer layers, each with 16 attention heads and a hidden dimension of 1280. The decoder is a standard transformer decoder with $L=6$ layers, consistent with the design in HSMR [xia2025hsmr]. We train SKEL-CF using the AdamW [loshchilov2017decoupled] optimizer with $\beta_{1}$, $\beta_{2}$ setting to 0.9, 0.999, weight decay setting to $1\times 10^{-4}$, a batch size of 64 and a learning rate of $1\times 10^{-5}$, preceded by a one-epoch warm-up. Training is conducted for 30 epochs on 8 NVIDIA A100 GPUs, taking approximately 120 hours in total. The hyper-parameters are set as $\lambda_{kp}=0.05$, $\lambda_{\beta}=0.0005$, $\lambda_{\theta}=0.001$, and $\lambda_{ref}=0.1$. Notably, our training schedule is significantly shorter than that of HSMR [xia2025hsmr], which trains for 100 epochs.
@@ 4 Experiment
Method 3DPW [3dpw] EMDB [kaufmann2023emdb] SPEC-SYN [spec-syn] MOYO [moyo] MOYO-HARD
@@
SMPL [loper2015smpl]-based Approaches
SPEC [spec-syn] 96.5 53.2 118.5 138.9 87.7 161.3 83.5 56.9 98.9 - - - - - -
CLIFF [li2022cliff] 69.0 43.0 81.2 103.5 68.3 123.7 128.5 55.8 139.0 154.6 109.3 155.7 - - -
HMR2.0a [goel20234dhuman] 69.8 44.4 82.2 97.8 61.5 120.0 133.3 55.8 153.0 - - - - - -
TokenHMR [dwivedi2024tokenhmr] 70.5 43.8 86.0 88.1 49.8 104.2 110.5 51.8 127.6 - - - - - -
WHAM [shin2024wham] 57.8 35.9 68.7 79.7 50.4 94.4 - - - - - - - - -
ReFit [wang2023refit] 57.6 38.2 67.6 91.7 55.5 106.2 103.6 51.3 116.3 - - - - - -
CLIFF [li2022cliff] 72.0 46.6 85.0 97.1 61.3 113.2 109.9 55.6 124.6 - - - - - -
HMR2.0b [goel20234dhuman] 81.3 54.3 93.1 118.5 79.2 140.6 150.7 67.6 172.9 123.3 90.4 142.2 - - -
CameraHMR [patel2024camerahmr] 62.7 38.7 73.4 73.2 43.9 85.6 66.0 37.0 79.1 85.1 50.2 95.3 96.0 61.9 111.8
SKEL [SKEL]-based Approaches
HSMR [xia2025hsmr] 81.5 54.8 - - - - - - - 104.5 79.6 120.1 120.0 97.7 140.5
@@
Table 2: Quantitative comparison of mesh recovery using the SMPL [loper2015smpl] human model on 3DPW [3dpw], EMDB [kaufmann2023emdb], SPEC-SYN [spec-syn], MOYO [moyo], and the more challenging MOYO-Hard dataset. The proposed SKEL-based method, SKEL-CF, achieves performance competitive with CameraHMR [patel2024camerahmr], with more pronounced improvements on complex motion scenarios as MOYO-Hard.
Figure 4: Visual comparison between the proposed SKEL-CF and CameraHMR [patel2024camerahmr]. The proposed SKEL-CF is built upon the SKEL [SKEL] representation, which enforces anatomically consistent joint motion, resulting in more natural pose predictions compared to the SMPL [loper2015smpl]-based CameraHMR [patel2024camerahmr] (best viewed in zoom). Additional visual examples are provided in the supplementary material.
@@ 4.1 Datasets and evaluation metrics
We evaluate SKEL-CF on five representative datasets covering diverse environments and motion complexities, following standard practice [patel2024camerahmr, xia2025hsmr, kanazawa2018hmr, dwivedi2024tokenhmr]. 3DPW [3dpw] provides in-the-wild videos with accurate 3D annotations captured by moving cameras. Human3.6M [human36m] includes controlled indoor actions (e.g., sitting, walking, greeting) with motion-capture ground truth. EMDB [kaufmann2023emdb] offers high-quality pose and shape sequences captured with electromagnetic sensors and handheld cameras. SPEC-SYN [spec-syn] is synthetic with explicit camera intrinsics and diverse perspectives for controlled camera evaluation. In contrast, MOYO [moyo] contains large-scale yoga sequences with extreme poses, frequent self-occlusions, and ground contact, providing the most challenging and diverse test scenario. Since most yoga videos in MOYO starts and ends with the same static poses, we select the front-view camera among the eight available viewpoints and remove the first and last 25% of frames to retain the complex and diverse motion segments, forming the curated MOYO-HARD subset. More details are provided in Appendix, Figure S5., where we additionally include visualizations of the removed first and last 25% segments of the MOYO videos.
@@
We compare to SKEL-based [SKEL] approaches in Table 1. Our proposed SKEL-CF achieves consistent and significant improvements over both HSMR [xia2025hsmr] and the optimization-based SKEL [SKEL] baseline across all benchmarks. SKEL-CF achieces significant improvement compared to previous approach arcoss all the three datasets. Notably, on the challenging MOYO [moyo] dataset, which involves large pose variations, self-occlusion, and ground contact, posing substantial challenges for accurate pose and shape estimation. SKEL-CF shows a clear advantage in this setting, achieving up to 18.6% in terms of MPJPE and 35.4% improvement on PA-MPJPE.
@@
We provide a visual comparison with HSMR [xia2025hsmr] in Figure 3. As shown, the proposed SKEL-CF achieves more accurate skeletal estimation and mesh reconstruction. More visual examples similar to Figure 3 are included in in Appendix, Figure S2.
Figure 5: Illustration of the mesh refinement process. The light green meshes represent the initial coarse estimations, while the dark green meshes denote the progressively refined final results.
@@ 4.3 Comparison with SMPL-based approaches
We further compare SKEL-CF with recent state-of-the-art SMPL-based human mesh recovery methods across four representative datasets: 3DPW [3dpw], EMDB [kaufmann2023emdb], SPEC-SYN [spec-syn], MOYO [moyo] and more challenging MOYO-HARD, as summarized in Table 2. Note that SMPL [loper2015smpl] is less constrained compared to SKEL [SKEL], enable achieving strong quantitative performance but occasionally producing anatomical inconsistency or unnatural joint motions(see Figure 1). Despite this inherent difference, SKEL-CF achieves comparable performance to CameraHMR [patel2024camerahmr], the strongest SMPL-based model trained on 4DHuman, demonstrating that kinematically constrained SKEL representations can reach a similar level of numerical precision while maintaining stronger physical plausibility.
Notably, on the challenging MOYO [moyo] dataset, which contains complex yoga sequences with strong self-occlusion and body–ground interactions, SKEL-CF achieves state-of-the-art performance in dense mesh prediction (PVE) and overall reconstruction accuracy. Furthermore, on the MOYO-HARD subset, focused on dynamic and highly articulated poses, SKEL-CF consistently outperforms all SMPL-based baselines across all metrics (MPJPE, PA-MPJPE, and PVE). This highlights its superior ability to maintain accurate and physically realistic motion representations under extreme articulations. These results demonstrate that SKEL-CF provides a scalable and anatomically faithful framework for human motion analysis, bridging the gap between computer vision and biomechanics. For completeness, we also report the 2D keypoint results in Table 3. Our approach achieves comparable accuracy under PCK@0.1, indicating that the projected mesh provides a consistent and accurate fit to the 2D image plane.
@@
We provide a visual comparison with the SMPL [loper2015smpl]-based state-of-the-art method CameraHMR [patel2024camerahmr] in Figure 4. As shown, the proposed SKEL-CF produces anatomically more consistent joint motions, benefiting from the structural constraints of the SKEL [SKEL] model. In contrast, CameraHMR tends to generate unnatural joint configurations under complex motion scenarios. Additional visualizations similar to Figure 4 are provided in Appendix, Figure S1.

### v4

@@ 3.2 SKEL-CF
Methods 3DPW [35] Human3.6M [9] MOYO [33] MOYO-HARD
MPJPE $\downarrow$ PA-MPJPE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$
Two-stage Approaches
HMR2.0 + SKEL fit [39, 6] 81.0 54.4 53.6 34.1 130.5 93.7 - -
CameraHMR + SKEL fit [39, 27] 70.4 41.8 - - 75.5 49.9 88.7 61.4
End to End Approaches
HSMR [39] 81.5 54.8 50.4 32.9 104.5 79.6 120.0 97.7
SKEL-CF (Ours) 61.5 38.7 39.0 31.2 85.0 51.4 90.0 61.5
Table 1: Quantitative comparison of skeleton and surface mesh recovery using the SKEL [15] human model on 3DPW [35], Human3.6M [9], and MOYO [33]. The proposed SKEL-CF delivers consistent accuracy gains over previous end-to-end approach. denotes results reported by [39], denotes fitting the outputs of [27] into the SKEL [15] model.
@@
Given an input image, the encoder first encodes it into visual tokens and, together with camera intrinsic information, extracts contextual features through standard self-attention.
Using these features, the encoder produces initial estimates $\boldsymbol{\Theta}_{0}=\{\boldsymbol{\theta}_{0},\boldsymbol{\beta}_{0},\boldsymbol{\pi}_{0}\}$. These initial predictions are then progressively refined through $T$ decoder layers using cross-attention with the encoded visual features as context. The final output of SKEL-CF is $\boldsymbol{\Theta}_{T}=\{\boldsymbol{\theta}_{T},\boldsymbol{\beta}_{T},\boldsymbol{\pi}_{T}\}$.
SKEL learning objectives.
We learn the target parameters $\boldsymbol{\hat{\Theta}}=\{\boldsymbol{\hat{\theta}},\boldsymbol{\hat{\beta}},\boldsymbol{\hat{\pi}}\}$ using a loss integrating keypoint-level and parameter-level supervision:
$\mathcal{L}_{\text{skel}}=\underbrace{\lambda_{kp}\mathcal{L}_{\text{kp}}}_{\text{keypoint-level}}+\underbrace{\lambda_{\beta}\mathcal{L}_{\beta}+\lambda_{\theta}\mathcal{L}_{\theta}}_{\text{parameter-level}}.$ (1)
where $\lambda_{kp}$, $\lambda_{\beta}$, and $\lambda_{\theta}$ are hyperparameters controlling the relative weight of each term. This formulation allows us to simultaneously supervise the reconstructed 3D/2D joints and the underlying pose and shape parameters. Note that no direct camera extrinsic supervision $\boldsymbol{\hat{\pi}}$ is provided; the ability to predict camera extrinsics is only implicitly learned through weak 2D keypoint reprojection, making the training more challenging.
Keypoint-level supervision.
Given the pose $\boldsymbol{\theta}$ and shape $\boldsymbol{\beta}$, we first reconstruct 3D joint positions $\mathbf{J}_{3d}\in\mathbb{R}^{K\times 3}$ via forward kinematics, where $K$ is the number of joints. These 3D joints are projected onto the image plane using the predicted camera parameters to obtain 2D joint locations $\mathbf{J}_{2d}\in\mathbb{R}^{K\times 2}$. The coarse keypoint loss $\mathcal{L}_{\text{kp}}$ enforces consistency with pseudo-ground-truth 3D and 2D joints:
@@
Parameter-level supervision.
Since the proposed dataset 4DHuman-SKEL provides reliable pose $\hat{\boldsymbol{\theta}}$ and shape $\hat{\boldsymbol{\beta}}$ parameters, we impose supervision directly on these values:
$\mathcal{L}_{\beta}=\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{1},\quad\mathcal{L}_{\theta}=\|\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}\|_{1}.$ (3)
Figure 3: Visual comparison between the proposed SKEL-CF and HSMR [39]. Our proposed SKEL-CF achieves more precise skeletal estimations (best viewed in zoom). Additional visual results are provided in the supplementary material.
@@
SKEL-CF adopts a coarse-to-fine refinement strategy. The encoder first produces an initial estimate of the parameters, denoted as $\boldsymbol{\Theta}_{0}=\{\boldsymbol{\theta}_{0},\boldsymbol{\beta}_{0},\boldsymbol{\pi}_{0}\}$. The decoder then refines these predictions across $L$ layers, where the $i$-th layer outputs $\boldsymbol{\Theta}_{i}=\{\boldsymbol{\theta}_{i},\boldsymbol{\beta}_{i},\boldsymbol{\pi}_{i}\}$. The final layer produces the refined prediction $\boldsymbol{\Theta}_{L}$. Both the encoder output and the final decoder output are optimized using the same objective function $\mathcal{L}_{\text{skel}}$:
@@
Iterative refinement.
To stabilize training, we adopt an iterative refinement strategy with layer-wise supervision. Specifically, each decoder layer predicts a residual update to the parameters based on the output of the previous layer, progressively refining the estimate. To facilitate effective learning at each stage, we apply auxiliary supervision [44, 3] to every decoder layer using the same refinement loss.
@@
where $L$ denotes the total number of decoder layers, and $\hat{\boldsymbol{\theta}}_{i}$ denotes the corresponding ground-truth pose parameters. We observe that optimizing $\mathcal{L}_{\text{skel}}$ provides similar performance. Since computing 2D keypoint predictions via the SKEL forward process is computationally expensive, we restrict optimization to the pose parameters.
Note that coarse-to-fine strategy, auxiliary supervision, and iterative refinement are widely adopted in the DETR-based object detection frameworks [44, 3]. The experiments to demonstrate the effectiveness of iterative refinement are provided in Appendix S2.
@@
The overall loss function combines the above components for both the coarse encoder prediction and the iterative refinement in decoder:
@@
Following HSMR [39], we adopt ViTPose-H [41], pretrained on COCO [22], as the encoder backbone. The hyper-parameters are set as $\lambda_{kp}=0.05$, $\lambda_{\beta}=0.0005$, $\lambda_{\theta}=0.001$, and $\lambda_{ref}=0.1$. More implementation details are provided in Appendix S3.
@@ 4 Experiment
Method 3DPW [35] EMDB [14] SPEC-SYN [17] MOYO [33] MOYO-HARD
@@
SMPL [23]-based Approaches
SPEC [17] 96.5 53.2 118.5 138.9 87.7 161.3 83.5 56.9 98.9 - - - - - -
CLIFF [21] 69.0 43.0 81.2 103.5 68.3 123.7 128.5 55.8 139.0 154.6 109.3 155.7 - - -
HMR2.0a [6] 69.8 44.4 82.2 97.8 61.5 120.0 133.3 55.8 153.0 - - - - - -
TokenHMR [4] 70.5 43.8 86.0 88.1 49.8 104.2 110.5 51.8 127.6 - - - - - -
WHAM [30] 57.8 35.9 68.7 79.7 50.4 94.4 - - - - - - - - -
ReFit [36] 57.6 38.2 67.6 91.7 55.5 106.2 103.6 51.3 116.3 - - - - - -
CLIFF [21] 72.0 46.6 85.0 97.1 61.3 113.2 109.9 55.6 124.6 - - - - - -
HMR2.0b [6] 81.3 54.3 93.1 118.5 79.2 140.6 150.7 67.6 172.9 123.3 90.4 142.2 - - -
CameraHMR [27] 62.7 38.7 73.4 73.2 43.9 85.6 66.0 37.0 79.1 85.1 50.2 95.3 96.0 61.9 111.8
SKEL [15]-based Approaches
HSMR [39] 81.5 54.8 - - - - - - - 104.5 79.6 120.1 120.0 97.7 140.5
@@
Table 2: Quantitative comparison of mesh recovery using the SMPL [23] human model on 3DPW [35], EMDB [14], SPEC-SYN [17], MOYO [33], and the more challenging MOYO-HARD dataset. The proposed SKEL-based method, SKEL-CF, achieves performance competitive with CameraHMR [27], with more pronounced improvements on complex motion scenarios as MOYO-HARD.
Figure 4: Visual comparison between the proposed SKEL-CF and CameraHMR [27]. The proposed SKEL-CF is built upon the SKEL [15] representation, which enforces anatomically consistent joint motion, resulting in more natural pose predictions compared to the SMPL [23]-based CameraHMR [27] (best viewed in zoom). Additional visual examples are provided in the supplementary material.
@@ 4.1 Datasets and evaluation metrics
We evaluate SKEL-CF on five representative datasets covering diverse environments and motion complexities, following standard practice [27, 39, 11, 4]. 3DPW [35] provides in-the-wild videos with accurate 3D annotations captured by moving cameras. Human3.6M [9] includes controlled indoor actions (e.g., sitting, walking, greeting) with motion-capture ground truth. EMDB [14] offers high-quality pose and shape sequences captured with electromagnetic sensors and handheld cameras. SPEC-SYN [17] is synthetic with explicit camera intrinsics and diverse perspectives for controlled camera evaluation. In contrast, MOYO [33] contains large-scale yoga sequences with extreme poses, frequent self-occlusions, and ground contact, providing the most challenging and diverse test scenario. Since most yoga videos in MOYO starts and ends with the same static poses, we select the front-view camera among the eight available viewpoints and remove the first and last 25% of frames to retain the complex and diverse motion segments, forming the curated MOYO-HARD subset. More details are provided in Appendix, Figure S1., where we additionally include visualizations of the removed first and last 25% segments of the MOYO videos.
@@
We compare SKEL-based methods in Table 1. Among end-to-end approaches, SKEL-CF consistently outperforms prior work HSMR [39] across all datasets. On the challenging MOYO [33] benchmark, SKEL-CF achieves up to 18.6% MPJPE and 35.4% PA-MPJPE improvements, demonstrating strong robustness under large pose variations and occlusions. We also evaluate two-stage baseline, HMR2.0 + SKEL fit and CameraHMR + SKEL fit. While CameraHMR + SKEL fit achieves competitive results on MOYO and MOYO-HARD, its performance drops on 3DPW, indicating limited stability across datasets. Moreover, the fitting-based approach is computationally expensive (approximately 4 min/img on RTX 3090), whereas SKEL-CF runs in a single forward pass (approximately 0.8 s/img), making it both more efficient and more reliable.
@@
We provide a visual comparison with HSMR [39] in Figure 3. As shown, the proposed SKEL-CF achieves more accurate skeletal estimation and mesh reconstruction. More visual examples similar to Figure 3 are included in Appendix, Figure S3.
Figure 5: Illustration of the mesh refinement. The light green meshes represent the initial coarse estimations, while the dark green meshes denote the progressively refined final results.
@@ 4.3 Comparison with SMPL-based approaches
We further compare SKEL-CF with recent state-of-the-art SMPL-based human mesh recovery methods across four representative datasets: 3DPW [35], EMDB [14], SPEC-SYN [17], MOYO [33] and more challenging MOYO-HARD, as summarized in Table 2. Note that SMPL [23] is less constrained compared to SKEL [15], enable achieving strong quantitative performance but occasionally producing anatomical inconsistency or unnatural joint motions(see Figure 1). Despite this inherent difference, SKEL-CF achieves comparable performance to CameraHMR [27], the strongest SMPL-based model trained on 4DHuman, demonstrating that kinematically constrained SKEL representations can reach a similar level of numerical precision while maintaining stronger physical plausibility.
Notably, on the challenging MOYO [33] dataset, which contains complex yoga sequences with strong self-occlusion and body–ground interactions, SKEL-CF achieves state-of-the-art performance in dense mesh prediction (PVE) and overall reconstruction accuracy. Furthermore, on the MOYO-HARD subset, focused on dynamic and highly articulated poses, SKEL-CF consistently outperforms all SMPL-based baselines across all metrics (MPJPE, PA-MPJPE, and PVE). This highlights its superior ability to maintain accurate and physically realistic motion representations under extreme articulations. These results demonstrate that SKEL-CF provides a scalable and anatomically faithful framework for human motion analysis. For completeness, we also report the 2D keypoint results in Table 3. Our approach achieves comparable accuracy under PCK@0.1, indicating that the projected mesh provides a consistent and accurate fit to the 2D image plane.
@@
We provide a visual comparison with the SMPL [23]-based state-of-the-art method CameraHMR [27] in Figure 4. As shown, the proposed SKEL-CF produces anatomically more consistent joint motions, benefiting from the structural constraints of the SKEL [15] model. In contrast, CameraHMR tends to generate unnatural joint configurations under complex motion scenarios. Additional visualizations similar to Figure 4 are provided in Appendix, Figure S2.